how much would you buy/ sell capacity if the forecast 1671, inventory on hand is 39, production schedule 1882, production after adjustment 1864, 1st shift capacity is 1400.
Answer #1

Forecast (F) = 1671

Inventory on hand (I) = 39

Production after adjustment (P) - 1864

1st shift capacity (C) = 1671

buy/sell capacity = Max (F,P) - I - C = Max (1671,1864) - 39 - 1400 = 425

So, they must buy capacity pf 425 units